OpenStudy (anonymous):

let f(x)=2x-x^2...... find (f(3+h))-(f(3)/h)

hero (hero):

f(3 + h) = 2(3 + h) - (3 + h)^2 f(3) = 2(3) - 3^2

OpenStudy (anonymous):

ah thank you thank you!!